OpenStudy (anonymous):

How do i solve for z? (z+1)^2 = 3 + 4i

hartnn (hartnn):

write z = x+iy

hartnn (hartnn):

\((x+iy+1)^2 = ([x+1] + iy )^2 =.....\) use \( (a+b)^2 = a^2+2ab+b^2\) formula

OpenStudy (anonymous):

Oh.. let me try that out

OpenStudy (anonymous):

What I have so far now is... \[(x+1)^{2} +2iy(x+1) - y ^{2} = 3 +4i\]

hartnn (hartnn):

yes, correct. go on

OpenStudy (anonymous):

Thanks. I will try and expand and see what happens

OpenStudy (anonymous):

Hmm.. not sure if I went the wrong way here... \[x ^{2} +2x - y ^{2} +1 + i(2xy +2y) = 3 + 4i\] Not sure how to go about putting in z = x + iy

hartnn (hartnn):

actually, i can think of a better way... why not let u = z+1 then u^2 = 3+4i

hartnn (hartnn):

it would be much easier to find u then we can always re substitute u = z+1

hartnn (hartnn):

and letting u = x+iy

OpenStudy (anonymous):

I see. And I believe I can use some equations I have here letting a = 3 and b =4 Thanks! I am going to try that approach

OpenStudy (anonymous):

Ok, I got close it seems.. \[z = 1 +\frac{ 1 }{ \sqrt{2} }i and z = -3 - \frac{ 1 }{ \sqrt{2} }i\] The solutions are suppose to be z = 1+i and z =-3 -i

hartnn (hartnn):

x^2 - y^2 = 3 2xy = 4 you got these ?

OpenStudy (anonymous):

No.. I used some formulas that used alpha and beta.. Probably should have worked to but I see you replaced I see how you got those though... Let me try 1 more time!! Thanks

terenzreignz (terenzreignz):

I have an idea, but you guys would probably curse me for even suggesting it ^_^ ...polar form, anyone?

hartnn (hartnn):

haha, why not! if it reduces any efforts :P

terenzreignz (terenzreignz):

\[\Large 3+4\color{blue}i = 5\left[\cos\theta + \color{blue}i\sin\theta\right]\]

terenzreignz (terenzreignz):

So that... to get the solution of \[\Large (z+1)^2 = 3+4\color{blue}i = 5\left[\cos\theta + \color{blue}i\sin\theta\right]\]

OpenStudy (anonymous):

I going to try both as this question is before the section on polar form but I need to know both

terenzreignz (terenzreignz):

We just have to get the two square roots of the complex-number to the right....

terenzreignz (terenzreignz):

Before polar form, huh? Should I proceed anyway? ^_^

hartnn (hartnn):

yes, we'll discuss both methods here.

OpenStudy (anonymous):

Yes. Proceed haha

terenzreignz (terenzreignz):

All right. Let me just get this off the table: ^_^ \[\Large z = r\left[\cos\theta + \color{blue}i\sin\theta\right]\]\[\Large z^\color{green}p=r^\color{green}p\left[\cos \color{green}p\theta + \color{blue}i\sin \color{green}p\theta \right]\] This is DeMoivre's theorem... it works for fractional powers too... (like square roots).... sort of.... So...

terenzreignz (terenzreignz):

\[\Large (z+1)^2 = 5\left[\cos\theta + \color{blue}i\sin\theta\right]\] Get the square root of both sides, we get: \[\Large z+1 =\sqrt 5\left[\cos\frac{\theta}2 + \color{blue}i\sin\frac{\theta}2\right]\] Since square root is the same as raising to a power 1/2.

OpenStudy (anonymous):

Relating to first method...Using x^2 - y^2 = 3 2xy = 4 I got.... let y = 2/x and plugged to top equation and got for x values... x = +-2 +-i i got the corresponding y's although not sure what to do now

hartnn (hartnn):

NOTE : in u = x+iy x and y are real numbers!

OpenStudy (anonymous):

The polar form is looking good. I have seen those equations but needed to check how theta got divided by 2 but I see now, like you said

terenzreignz (terenzreignz):

We don't (actually) know what \(\theta\) is... (ugly, whether in degrees or radians) But we DO know that \[\Large \cos \theta = \frac35\] and this is actually all we need ^_^ Remembering half-angle identities... \[\Large \cos\frac\theta 2=\sqrt{\frac{1+\cos\theta}{2}}=\sqrt{\frac4{5}}\]after substituting. similarly...\[\Large \sin\frac\theta 2=\sqrt{\frac{1-\cos\theta}{2}}=\sqrt{\frac1{5}}\]

OpenStudy (anonymous):

Alright I will use the real numbers for x and y

terenzreignz (terenzreignz):

Back to the polar form, we get: \[\Large z+1 =\sqrt 5\left[\cos\frac{\theta}2 + \color{blue}i\sin\frac{\theta}2\right]\]\[\Large z+1 =\sqrt 5\left[\sqrt{\frac45}+ \color{blue}i\sqrt{\frac15}\right]=2+\color{blue}i\] And the rest is history ^_^

OpenStudy (anonymous):

Nice! And I believe I can find the other solution

terenzreignz (terenzreignz):

As for the other root, we could go back to here: \[\Large (z+1)^2 = 5\left[\cos\theta + \color{blue}i\sin\theta\right]\] And just note that \[\Large (z+1)^2 = 5\left[\cos(\theta+2\pi) + \color{blue}i\sin(\theta+2\pi)\right]\] And again, get the square root... \[\Large z+1 = \sqrt5\left[\cos\frac{(\theta+2\pi)}2 + \color{blue}i\sin\frac{(\theta+2\pi)}2\right]\] \[\Large z+1 =\sqrt 5\left[\cos\left(\frac{\theta}2+\pi\right) + \color{blue}i\sin\left(\frac{\theta}2+\pi\right) \right]\]

terenzreignz (terenzreignz):

Oh sorry, feel free to proceed, then ^_^

OpenStudy (anonymous):

Thank you both!! I just finished using the other method and just to be safe.. We using half-angle identities for this solution as well?

terenzreignz (terenzreignz):

Of course, and of course, the fact that... \[\Large \cos(\alpha+\pi) = -\cos(\alpha)\\\Large \sin(\alpha+\pi) = -\sin(\alpha)\] good hunting ^_^ Note that I didn't include the +/- bit in the half-angle identity because the argument \(\theta\) is in the first quadrant (due to the real and imaginary parts of 3+4i both being positive)

OpenStudy (anonymous):

Great! All I do is change the sign thanks to pie

terenzreignz (terenzreignz):

Yup. And you get the other answer, am I right? ^_^

OpenStudy (anonymous):

Yep!! I just did. Thanks a bunch!